OpenStudy (anonymous):

simplify: √ 2 (x- √ 5) Help please :(

OpenStudy (anonymous):

\[\sqrt{2x} - \sqrt{10}\]

OpenStudy (anonymous):

@antoni7 thank you :)

OpenStudy (anonymous):

If you were to find the Roots it would be like that : √ 2 x- √ 10 = 0 x = √ 10 / √ 2 x = √ 5

OpenStudy (anonymous):

no problem (: